### v2

# Inferring Compositional 4D Scenes without Ever Seeing One

Abstract
Scenes in the real world are often composed of several static and dynamic objects. Capturing their 4-dimensional structures, composition and spatio-temporal configuration in-the-wild, though extremely interesting, is equally hard. Therefore, existing works often focus on one object at a time, while relying on some category-specific parametric shape model for dynamic objects. This can lead to inconsistent scene configurations, in addition to being limited to the modeled object categories. We propose COM4D (Compositional 4D), a method that consistently and jointly predicts the structure and spatio-temporal configuration of 4D/3D objects using only static multi-object or dynamic single object supervision. We achieve this by a carefully designed training of spatial and temporal attentions on 2D video input. The training is disentangled into learning from object compositions on the one hand, and single object dynamics throughout the video on the other, thus completely avoiding reliance on 4D compositional training data. At inference time, our proposed attention mixing mechanism combines these independently learned attentions, without requiring any 4D composition examples. By alternating between spatial and temporal reasoning, COM4D reconstructs complete and persistent 4D scenes with multiple interacting objects directly from monocular videos. Furthermore, COM4D provides state-of-the-art results in existing separate problems of 4D object and composed 3D reconstruction despite being purely data-driven. Code is available at https://github.com/insait-institute/COM4D.
Figure 1: Given a single video (bottom), our method reconstructs the entire 3D scene along with the individual dynamic objects (A), while maintaining spatial and temporal consistency through spatio-temporal attention mixing (C). The silhouettes (purple for human and orange for dog) correspond to the beginning of dynamic sequences. Our user study (for spatial correctness and temporal coherence) shows that the reconstructions obtained using the proposed attention mixing mechanism are clearly preferred over the baseline without mixing (B).
1 Introduction
Real scenes consist of multiple static and dynamic objects whose structures, compositional relationships, and spatio-temporal configurations evolve continuously over time. Capturing these factors jointly, without restrictive assumptions, requires solving reconstruction, decomposition, and temporal reasoning simultaneously. Existing approaches often sidestep this by focusing on a single object at a time or relying on category-specific parametric models for dynamic entities, which may result in inconsistent scene geometry, limited generalization, and strong inefficiencies whenever real-world objects or motions deviate from modeled priors. Consequently, despite its importance, object-decomposed 4D scene modeling in-the-wild remains highly challenging.
Single-object 3D from an image [49, 43, 98, 93, 96] has advanced rapidly, thanks to the large scale object [15], and powerful generative pipelines [25, 16, 48, 50] combining self-supervised visual priors [59, 79, 17], Variational Autoencoder (VAE) shape spaces [92, 85, 43, 89], and diffusion transformers [64]. Object-agnostic point-based 3D/4D scene reconstruction has progressed through self-supervised image priors, predicting the joint structure [82, 40, 34, 67, 81] or structure with parametric 4D fitting [12]. These successes hinge on simplifying assumptions: objects are static, or of known parametric class, data scale is massive, and structure is globally consistent.
Similarly restrictive, most approaches to dynamic 4D reconstruction are limited to a single deforming object [10, 94, 68, 65, 7], multiple objects [30, 12] defined by category-specific parametric models [53, 63], or are object-unaware  [81, 87, 83, 95, 11]. Typically, motion in training data is captured in controlled environments using active sensors, dedicated motion-tracking rigs, or via physical deformation models or carefully designed synthetic assets. Yet, real-world scenes violate all of these conveniences: multiple static and dynamic objects coexist, interact, occlude one another, and exhibit heterogeneous geometric and motion patterns that defy any unified prior. Consequently, as soon as objects move behind others, exhibit complex interactions, or undergo significant viewpoint changes, 4D structure can no longer be captured in many approaches, resulting in fragile representations. In other words, most 4D scene recovery often struggle with both consistency and persistence of objects. For the same reasons, large-scale, in-the-wild 4D multi-object data [32] is extremely scarce, making learning severely under-constrained. As a result, progress on multi-object 4D scene reconstruction has lagged far behind that of simpler settings. To the best of our knowledge, no existing model can infer a complete and persistent 4D representation of multiple static and dynamic objects in-the-wild using only monocular videos, without test-time optimization.
To break through this barrier, we pursue a fundamentally different approach. We show that the required spatio-temporal reasoning can be learned in the form of attentions. We do so separately from two sources that are easy to obtain: static multi-object observations for spatial structure, and single-object animations for temporal dynamics. Thus, we introduce COM4D, a compositional 4D reconstruction method that unifies these independently learned attentions at inference time, guided by a simple but powerful physical assumption: at every time instant, all scene elements are momentarily static, and their dynamics unfold by propagating object states forward in time. By iteratively alternating spatial and temporal reasoning — a process we call attention mixing (as illustrated in Figure Inferring Compositional 4D Scenes without Ever Seeing One.C), COM4D implicitly recovers the complete and persistent 4D structure of multiple interacting objects without ever seeing a single example of such data during training. Realizing this intuition required a series of careful design choices in representation, architecture, supervision, and inference, which collectively allow us to address a problem long considered exceptionally hard. We summarize our main contributions as follows:
  • We introduce attention parsing, a simple yet effective strategy that disentangles the learning of spatial and temporal reasoning from separate, complementary data sources, without compromising the quality of either.
  • We propose attention mixing, which unifies these independently learned attentions at inference time on an input of a video of any length, to achieve compositional 4D scene reconstruction. Thus, our model recovers multi-object, spatio-temporal structure despite never being explicitly trained on such data.
  • We show that this unified attention framework generalizes across diverse in-the-wild scenes, achieving coherent and persistent 4D reconstructions of multiple interacting objects, significantly outperforming specialized dynamic single-object or static-scene baselines.
2 Related Works
4D Object Reconstruction.
In earlier works, termed as non-rigid structure-from-motion, 4D reconstruction approaches primarily relied on low-rank assumptions [6, 75, 60, 14, 57] and physics-based priors [69, 61, 1]. Due to more practical results, subsequent works favored category-specific approaches [99, 23, 21, 38, 22, 62] using parametric human shape models [53, 63]. In order to handle any shape, methods [31, 2] later opted the score-distillation [66, 84, 26]. Considering both speed and generalizability to any shape, generative or diffusion-based approaches [46, 10, 94, 68] have become popular, as they have the potential to capture any shape persistently (including the occluded regions), while avoiding expensive test time adaptations. Noteably, [68] proposed to learn temporal self-attention in a local-global manner similar to [86] instead of borrowing the pretrained priors from video diffusion models [4, 20].
Composed 4D Scene Reconstruction.
The steady advancement of 3D scene reconstruction [76, 56, 73, 70, 40, 82, 79] have enabled learning-based approaches to unify static-dynamic scene 4D reconstruction [97, 28, 95, 81, 12, 18]. This has been largely possible through the novel unified appearance-3D representations [55, 35]. Still, most previous works target human motion [23, 72, 74, 51, 12, 71] or do not tackle the persistent object decomposed reconstruction [80, 39, 81, 11, 83, 91, 37]. Alternately, many current approaches [44, 45, 52, 39, 80] for 4D scenes use an online optimization paradigm limiting their efficiency. Specifically [13] reconstructs static-dynamic objects in their respective coordinates and recomposes them together with their poses and depth, through test-time optimization. Test-time optimization is also favored by other 4D complete scene reconstruction approaches [80, 39, 11].
Generative Scene Reconstruction.
Despite novel representations [55, 35] unlocking new capabilities [90, 27], many real-world applications still favor persistent one-mesh-per-object explicit geometry [98, 96]. Thanks to large scale data [15, 88], incredible advancements in fast single object reconstruction [49, 92, 43, 93, 89, 96, 98] have been made. However, surprisingly few works have tried integrating generative approaches for multi-object scene reconstruction, despite their potential in persistent, complete and object-aware results. Among the few, MIDI [29] proposes multi-instance attention in order to learn relative placement of objects from the object-wise masked conditional image embeddings. A similar approach with multi-instance attention and object-wise mask inputs is followed by [54]. Closest to our approach, the recent PartCrafter [47] circumvents object-wise masks, and instead denoises multiple latents by alternating in-part and inter-part attention with the part embeddings for object localization. Nevertheless, to the best of our knowledge, previous test-time optimization-free approaches (diffusion or single-step) do not solve object decomposed 4D reconstruction of scenes from video.
3 Method
Problem.
Consider a 4D scene composed of $N$ static (indexed by $i$) and $M$ dynamic (indexed by $j$) objects, recorded in a fixed camera monocular video over $F$ frames (indexed by $f$). We represent the images by their DINOv2 [59] embeddings $\{\smash{{}^{f}}\!\mathbf{y}^{j}\}$, with the static objects separated in the embedding $\mathbf{y}$. The goal of compositional 4D is to reconstruct the static object geometry latents $\mathcal{S}=\{\mathbf{z}^{i}\}$ and the dynamic ones $\mathcal{D}=\{\!{}^{f}\!\mathbf{z}^{j}\}$ conditioned on the image embeddings. Note that, we use the VAE latent quantity $z$ to also refer to an object geometry, for convenience.
Overview.
In this work, we aim to learn a compositional generative model for $\{\mathcal{S},\mathcal{D}\}$ without ever seeing such compositions. Consider a single Diffusion Transformer (DiT), $\mathbf{v}_{\theta}$ with parameters $\theta$, trained on the target distributions $p(\{\mathbf{z}^{i}\}|\mathbf{y})$: the static geometry composition distribution conditioned on the image embedding, and $p(\{{}^{f}\!\mathbf{z}|{}^{f}\!\mathbf{y}\})$: the dynamic single object shape distribution over each video frame, conditioned on each frame embedding. Given the trainings, we want $\mathbf{v}_{\theta}$ to also naturally model the target joint static-dynamic distribution $p(\mathcal{S},\mathcal{D}|\mathbf{y},\{{}^{f}\!\mathbf{y}^{j}\})$. Here, $\mathcal{D}$ may consist of multiple dynamic objects, unseen during the training of $\mathbf{v}_{\theta}$. In the following, we describe how we tackle the problem by designing $\theta$ and its corresponding learning.
The core of our method is a single DiT architecture that first learns to perform two distinct but complementary tasks: object aware reconstruction of 3D scenes and modeling the temporal dynamics of a deformable object. We achieve this through a dual-objective training strategy called Attention Parsing. At inference, a novel Attention Mixing mechanism allows us to combine these learned capabilities to conditionally generate complex 4D scenes with both static and dynamic components, a task for which the model was never explicitly trained. Finally, we enhance the temporal coherence of our generations by fine-tuning with Diffusion Forcing, cf. next section. Fig. 2 illustrates how Attention Parsing and Attention Mixing works.
Figure 2: Our attention parsing and mixing strategy. A single DiT model with shared weights is trained jointly on two datasets. (Top) During training with samples from DeformingThings [42], odd-indexed blocks perform multi-frame attention to capture temporal dynamics. (Bottom) When training with samples from 3D-FRONT [19], even-indexed blocks perform multi-instance attention to model spatial part decomposition. At inference, the same model applies an attention mixing mechanism. In each layer, spatial blocks (even-indexed) aggregate all latents from a single frame and process them jointly, conditioned on the full-scene image $y$ at that timestep. Temporal blocks (odd-indexed) then operate over all frames of each dynamic object separately, conditioned on their corresponding masked images. Masks are extracted from the video for each dynamic object using SAM [36], enabling temporally consistent object-specific reasoning.
3.1 Preliminaries
Diffusion Transformer for 3D. Our work builds upon TripoSG [43], a state-of-the-art image-to-mesh generative model. It consists of a shape Variational Autoencoder (VAE). The encoder of the VAE converts an object mesh into a latent feature which can then be decoded into the Signed Distance Field (SDF) through its decoder similar to [92]. A DiT [64, 43] conditioned on the DinoV2 [59] image embedding is then trained to denoise a noisy latent to the object shape’s VAE latent on Objaverse [15] and ShapeNet [8].
Diffusion Forcing. Different from the standard diffusion-based learning [48, 25], Diffusion Forcing [9] applies independent noise to the different latent vectors in the same data and denoises them together. The ability to process such mixed-noise inputs, where some latents are clean and others are noisy, is what makes this training scheme critical for our method’s requirements. First, it directly addresses the need for stable static guidance by training the model to handle conditioning on fully denoised static objects ($t=0$) while generating the remaining noisy dynamic latents. Second, it is essential for history-guided generation, as it enables a previously denoised frame f1f-1 to serve as a clean ($t=0$) context for generating the subsequent frame $f$. This ensures strong temporal consistency and coherent evolution of the dynamic objects.
3.2 Attention Parsing: Dual-Objective Training
Our key insight is to train a single DiT model such that it understands both spatial composition and dynamics by alternately training the model on two distinct datasets, with the transformer blocks performing different roles in each.
Dual-Dataset Strategy. At each training step, we sample with equal probability from either the 3D-FRONT [19] dataset, which provides static scenes with object-level decompositions, or the DeformingThings [42] dataset, which contains dynamic sequences of a single deforming object.
Alternating Block Roles. The DiT backbone consists of 21 transformer blocks. We assign complementary roles to these blocks depending on the data source. When training on a 3D-FRONT [19] sample, the even-indexed blocks are configured to perform multi-instance attention, enabling them to reason about the spatial relationships between different object parts in a scene. Conversely, when training on a DeformingThings [42] sample, the odd-indexed blocks are tasked with multi-frame attention, allowing them to capture the temporal dependencies across different frames of a sequence. The blocks not assigned to multi-instance attention (even blocks in DeformingThings [42]) or multi-frame attention (odd blocks in 3D-Front [19]) default to local self-attention, see Fig. 2.A.
Compositional Latent Space. Following our problem formulation, a scene is represented as a collection of $N+M$ latents (objects and/or frames). Each latent/token is a tensor $\mathbf{z}\in\mathbb{R}^{K\times C}$. To distinguish between the different objects/frames, we add a unique, learnable embedding in each token: an object embedding $e^{i}$ for 3D-FRONT samples and a frame embedding ${{}^{f}}\!e$ for DeformingThings samples. Similarly, a single frame embedding is added to the 3D-FRONT object latents and a single object embedding is added to the DeformingThings sample.
Diffusion Training Objective. Our architecture’s ability to reason about multiple components is enabled by the dual attention strategy. Local self-attention is applied independently to each latent’s tokens $\mathbf{z}$, while global reasoning is handled by transforming specific self-attention layers into multi-instance attention layers. The multi-instance attention allows the latent of each object $\mathbf{z}^{i}$ to attend to the rest $\{\mathbf{z}^{l}\}_{l=1}^{N}$ while updating itself. Similarly, the multi-frame attention each latent ${{}^{f}}\!\mathbf{z}$ using the rest ${{}^{F}_{l=1}}\!\{{{}^{l}}\!\mathbf{z}\}$. We write the multi-instance attention for multi-objects as below.
$\mathbf{z}^{i_{\text{out}}}=\text{Attention}(\mathbf{z}^{i},\{\mathbf{z}^{l}\}_{l=1}^{N}).$ (1)
The multi-frame attention follows similarly.
To train this architecture, we adapt the rectified flow objective for our dual-task. For the sake of brevity and correctness, we describe how the process works for static multiple objects in a scene. The rectified flow process for the multi-frame object latents follows similarly.
Crucially, to enable Diffusion Forcing [9], we sample an independent time step $t_{i}\!\in\![0,1]$ for each latent among $N$. Each clean latent $\mathbf{z}_{0}^{i}$ is then perturbed along its own linear trajectory by $\boldsymbol{\epsilon}^{i}\sim\mathcal{N}(0,\mathbf{I})$, a random Gaussian noise tensor:
$\mathbf{z}_{t_{i}}^{i}=t_{i}\mathbf{z}_{0}^{i}+(1-t_{i})\boldsymbol{\epsilon}^{i},$ (2)
The flow network $\mathbf{v}_{\theta}$ is trained to predict the velocity vector $\boldsymbol{\epsilon}^{i}-\mathbf{z}_{0}^{i}$ for each component based on its unique noisy state. The overall loss is the sum over all $N$ latents:
$\mathcal{L}_{S}=\mathbb{E}\left[\sum_{i=1}^{N}\left\|(\boldsymbol{\epsilon}^{i}-\mathbf{z}_{0}^{i})-\mathbf{v}_{\theta}(\mathbf{z}_{t_{i}}^{i},t_{i},\mathbf{y})\right\|^{2}\right].$ (3)
The conditioning tensor $\mathbf{y}$ is the background subtracted static object image embedding for eq. (3), where the samples are selected from the 3D-Front dataset [19]. During temporal training (on DeformingThings, see the equation in the supplementary), the conditioning tensor is ${{}^{f}}\!\mathbf{y}$, the unique image embedding corresponding to the $f$-th frame. The independent noising used in training the spatial and temporal denoising is vital, as it forces the model to handle latents at different stages of the denoising process, directly preparing it for history-guided generation at inference. The exact loss used for training is $\mathcal{L}_{S/T/R}$: static, temporal or the standard loss from TripoSG [43] for regularization.
3.3 Attention Mixing: Compositional 4D Denoising
As a consequence of Attention Parsing, our model can separate spatial and temporal reasoning. At the inference time, on a given compositional video input, the Attention Mixing strategy enables the joint denoising of complex 4D scenes of any combination of static and dynamic entities by modulating the information flow through the transformer blocks.
During a single denoising step at inference, the DiT blocks alternate their function to cohesively generate a 4D scene. First, the even-indexed spatial blocks reason about the global scene layout. To do this, they receive the concatenated latents of all static objects along with the latent representing the current frame of each dynamic object. This combined set forms a complete snapshot of the scene at a single moment. The block performs multi-instance attention across this entire set, using cross-attention keys and values derived from the single, global scene image to ensure all elements are placed correctly relative to one another. Subsequently, the odd-indexed temporal blocks model motion and deformation. These blocks process the latent sequence for each dynamic object separately, performing multi-frame attention over its history. The cross-attention keys and values for this operation are derived from the corresponding per-frame conditioning embeddings for that specific object, allowing the model to capture its unique temporal dynamics. Note that the strategy can handle any number of frames by sequentially propagating attention through a sliding temporal window (red block in Fig. 2 right) over all the frames, thus maintaining temporal consistency. Static object latents pass through these temporal blocks without being processed for motion. This flexible information routing allows the model to satisfy both the spatial constraints learned from 3D-FRONT [19] and the temporal dynamics learned from DeformingThings [42] within a single denoising pass. Algo. 1 describes Attention Mixing in a simple code.
Algorithm 1 Single Denoising Pass with Attention Mixing
# (N,M): Number of (Static, Dynamic) Objects
# y: Full scene condition; Z: Latent matrix ((N+M)xF)
# Y: Masked frame-wise conditions (MxF)
for block in transformer.blocks:
if is_spatial(block):
for f in 1..F:
       Z’[:,f] = block(Z[:,f],y,num_instances=N+M)
else:
for i in 1..N:
       Z’[i,:] = block(Z[i,:],y,num_instances=1)
for j in N..M:
       Z’[j,:]=block(Z[j,:],Y[j-N,:],num_instances=F)
    Z = Z’
return Z
Input
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Ours
Input
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Ours
Figure 3: Qualitative results on temporal sequences. Top rows show input frames; bottom rows show our generated reconstructions from two vertically stacked camera views. The examples shown are from content produced by ChatGPT [58] and animated with Wan [78], the CMU Panoptic dataset [32] (sequences 160401_ian3 and 160906_ian2), and the PROX dataset [24] (N3OpenArea_00158_02). Our method maintains temporal consistency and spatial realism across both real and synthetic sources.
3.4 Implementation Details
We build on the 21-block DiT backbone of TripoSG [43]. Even-indexed blocks operate in the spatial mode, while odd-indexed blocks operate in the temporal mode. This depth-wise alternation follows the multi-instance reasoning intuition [47] while keeping the original DiT [43] depth and width unchanged. We fully fine-tune all weights of the pretrained TripoSG DiT. We cap the supervision targets at 8 parts for the spatial path and 8 frames for the temporal. At each iteration, with probability $0.3$ we train on a monolithic sample (single part, single frame) to regularize the model and preserve its learned object prior ($\mathcal{L}_{R}$).
Training uses two stages on a single NVIDIA H200 GPU. In stage 1 (8k steps), we fine-tune the model with a batch size of $50$ and a learning rate of $1\times 10^{-4}$. In stage 2 (12k steps), we keep the batch size at $50$, enable Diffusion Forcing [9], and lower the learning rate to $1\times 10^{-5}$. The full training time is about 2 days.
4 Experiments
Test Datasets.
For 3D evaluation, we use the 3D-FRONT [19] test set following MIDI’s protocol [29]. For 4D evaluation, we construct two complementary sets: (i) an Objaverse-based subset starting from the animated object list released with Puppet-Master [41], from which we select 40 high-quality assets (clean geometry and faithful textures) from Objaverse [15]; and (ii) a DeformingThings4D [42] subset comprising 30 human and 30 animal animations. For every object or sequence, we render image sequences from a single fixed, calibrated camera and use the resulting frames as inputs for evaluation.
Evaluation Protocol.
We perform evaluation on three cases: Compositional 4D, 4D object and 3D scenes. For all methods we report runtime, fidelity, and task-specific structural quality. Runtimes are computed from the average inference time in NVIDIA H200.
For the static 3D scenes, we compare against MIDI-3D [29] and PartCrafter [47]. For a fair comparison, we provide MIDI-3D with ground-truth instance masks at inference, a step not required by PartCrafter or our method. For the 4D animation task, we include state-of-the-art generative approaches L4GM [68] and GVFD [94], the mesh-based V2M4 [10], and a frame-wise TripoSG [43] baseline to ablate the effects of temporal modeling. The token budget is fixed at 512 per part for all 3D methods and 1024 per frame for all mesh-based 4D methods.
Geometric fidelity is measured using the Chamfer Distance (CD) [3, 5] and F-Score [77] (at a 0.1 threshold), where lower CD and higher F-Score are better. For 3D scenes, we concatenate all parts into a single mesh before evaluation, whereas for 4D sequences, we compute the metrics per-frame and then average. To assess structural quality, we use two Intersection-over-Union (IoU) metrics based on a consistent $64^{3}$ voxelization. In 3D, we measure part independence via a pairwise IoU, where lower scores indicate better separation between parts. In 4D, we evaluate accuracy with a per-frame IoU against the ground truth, where higher scores are better. Note that for Gaussian-based baselines (L4GM, GVFD), we convert their outputs to point clouds for CD and F-Score evaluation, but omit the IoU metric for lack of reliable watertight meshes. All reported scores are the mean values over the test set.
4.1 Compositional 4D Reconstruction
Compositional 4D forms our primary experimental setup and is therefore our first task. We provide experimental results using various real and synthetic sequences. Fig. 3 provides the qualitative results on the synthetic, i.e., generated video on the top row and on the CMU Panoptic sequences [32] in the bottom row. Fig. 3 shows that our method completes the reconstruction capturing the inter-object interactions with surprising accuracy, thanks to the attention mixing strategy where dynamic objects can attend to its history as well as the static objects in the scene. We provide more qualitative/quantitative results for our method, with the impact of attention mixing in §4.4.
w/o Mixing w/ Mixing w/o Mixing w/ Mixing
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 4: Visualizations with and without our Attention Mixing strategy. Results are for 160401_ian3 at frame 1180 (starting frame: 1100) and 170915_office1 at frame 670 (starting frame: 590). Gray points denote ground truth.
4.2 Single Object 4D
Unlike the compositional case, several approaches exist for 4D object-aware generative reconstruction. We show qualitative comparisons in Fig. 5 in a novel view. We observe that our method captures the shape details accurately. TripoSG [43] provides strong static reconstruction, while it along with V2M4 [10], struggle on temporal consistency for certain frames. Similarly, GVFD [94] and L4GM [68] produce textured Gaussians that look plausible from the input view, but their underlying 3D geometry shows inconsistency when rendered from novel viewpoints.
Input Ours TripoSG V2M4 GVFD L4GM
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 5: Qualitative 4D generation comparisons for two subjects (top two rows: Ninja, bottom two rows: Amy) at two time steps. The first column shows the input frames, and subsequent columns show a fixed pose rendered view from each method. V2M4 fails in a few samples, e.g., the last row input.
The quantitative evaluations in Tab. 1 reflect the same observations. On DeformingThings [42], our approach significantly outperforms all baselines across every metric; for instance, our IoU of 0.4191 shows a substantial gain over the next best method. On Objaverse [15], our approach again achieves the best F-Score and IoU, indicating better shape completeness and overlap, while remaining highly competitive in Chamfer Distance against TripoSG [43]. Overall, these results validate the geometric accuracy and completeness of our method’s reconstructions on both moderate and large dynamics.
Table 1: Method comparison across datasets. Lower is better for CD; higher is better for F-Score and IoU. The best score is shown bold while the second best is shown italicized. Runtimes were averaged to reflect the reconstruction of a single frame.
DeformingThings [42] Objaverse [15] Runtime
Method CD $\downarrow$ F-Score $\uparrow$ IoU $\uparrow$ CD $\downarrow$ F-Score $\uparrow$ IoU $\uparrow$ Sec. $\downarrow$
V2M4 [10] 0.1678 0.4759 0.1533 0.1595 0.4903 0.2135 61.98
L4GM [68] 0.2633 0.3167 0.2308 0.3236 1.22
GVFD [94] 0.2806 0.2706 0.2728 0.2856 2.87
TripoSG [43] 0.1558 0.5179 0.1784 0.1107 0.6585 0.2874 2.59
Ours 0.1144 0.8388 0.4191 0.1205 0.7349 0.3413 4.48
4.3 3D Scene Reconstruction
Fig. 6 provides a visual comparison of our method against previous generative approaches: PartCrafter and MIDI, on the 3D-Front [19] examples. Our method reconstructs complete and detailed object layouts, being more faithful to the input. In contrast, PartCrafter [47] occasionally produces partial or low-quality meshes, while MIDI [29] struggles with occlusions.
Table 2: 3D Scene Generation on 3D-FRONT and 3D-FRONT-Occluded. Lower is better for CD and IoU; higher is better for F-Score. The best score is shown bold while the second best is shown italicized.
3D-FRONT [19] 3D-FRONT-Occluded [19] Runtime
Method CD $\downarrow$ F-Score $\uparrow$ IoU $\downarrow$ CD $\downarrow$ F-Score $\uparrow$ Seconds $\downarrow$
MIDI [29] 0.1445 0.7829 0.0027 0.1781 0.7387 3.51
PartCrafter [47] 0.1751 0.7569 0.0017 0.1951 0.7461 2.60
Ours 0.0909 0.8069 0.0018 0.1256 0.7521 2.63
The quantitative evaluation makes these even clearer. Tab. 2 compares our method with recent baselines on the 3D-FRONT dataset [19]. Our approach demonstrates state-of-the-art performance, achieving the best Chamfer Distance (0.0909) and F-Score (0.8069). This represents a significant improvement in geometric accuracy, surpassing the second-best method, MIDI [29]. This advantage holds even on the challenging 3D-FRONT-Occluded split.
Input Ours PartCrafter [47] MIDI [29]
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 6: Qualitative comparison across methods. Our approach generates more consistent and detailed structures compared to PartCrafter [47] and MIDI [29].
4.4 Ablation Study
We conduct a series of ablation studies to validate the effectiveness of our key architectural components: the use of distinct static/dynamic embeddings, the Diffusion Forcing training scheme, and our attention mixing strategy.
Quantitative Analysis. The quantitative results in  Tab. 3 validate our design choices against a baseline model. Adding the Static/Dynamic Embeddings yields the most significant performance gain; on DeformingThings, this drops the Chamfer Distance (CD) from 0.1525 to 0.1284 and nearly doubles IoU from 0.2018 to 0.4034, highlighting the importance of disentangling static and dynamic representations. Diffusion Forcing also provides a notable improvement, particularly for IoU, which enhances temporal consistency. Our full model COM4D, combining both components, achieves the best overall performance.
Analysis of Attention Mixing. Fig. 7 qualitatively validates our Attention Mixing strategy. Without it, the model fails to ground dynamic objects in their static context, resulting in artifacts where the cat overlaps with the lamp stand and humans are not sitting where they are supposed to. In contrast, by alternating between spatial and temporal attention in the Attention Mixing strategy, our model makes dynamic objects aware of static placements. This leads to reconstructions that are significantly more spatially accurate, validating our strategy for modeling plausible temporal evolutions in multi-object scenes.
We report quantitative results in two forms. Because no compositional 4D dataset with object-level decomposition exists, our main evaluation relies on user studies in Inferring Compositional 4D Scenes without Ever Seeing One.B, where our method is preferred over a standard generative baseline (without Attention Mixing) by a factor of 12 (87% vs. 6.9%). We also evaluate on CMU Panoptic [32, 33] point clouds by registering only the first frame reconstruction, where Attention Mixing reduces the average CD from 35.91 cm to 7.42 cm. This demonstrates that COM4D can accurately capture multiple dynamic shape evolutions in long sequences >90>90 frames, with surprisingly small drifts. Fig. 4 visualizes the reconstructions of ours and the baseline. Further details and additional quantitive analysis on synthetic compositional data appear in the supplementary material.
Table 3: Ablation study on 3D-FRONT and DeformingThings datasets. “Baseline” refers to the model without static/dynamic embeddings or diffusion forcing, while “Ours” uses both. Lower is better for CD and IoU in 3D-FRONT; higher is better for F-Score and IoU in DeformingThings.
3D-FRONT [19] DeformingThings [42]
Method CD $\downarrow$ F $\uparrow$ IoU $\downarrow$ CD $\downarrow$ F $\uparrow$ IoU $\uparrow$
Baseline 0.1668 0.6347 0.0025 0.1525 0.7854 0.2018
+ Static/Dynamic Emb. 0.1044 0.7888 0.0017 0.1284 0.9350 0.4034
+ Diffusion Forcing 0.1247 0.7152 0.0021 0.1488 0.8189 0.4271
Ours 0.0909 0.8069 0.0018 0.1144 0.8388 0.4191
Input w/o Mixing w/ Mixing
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 7: Qualitative comparisons on Attention Mixing. Mixing (right) provides far better static-dynamic compositions compared to without Mixing (middle).
5 Conclusion
In this work, we introduced COM4D, a novel approach for reconstructing compositional 4D scenes from monocular videos, without requiring direct supervision on such scarce data. Our work is motivated by the inherent difficulty of acquiring real world compositional 4D for training, which in fact, may not be resolved in the near future. We instead, address the problem by proposing a unique training strategy called Attention Parsing, which factorizes the learning of spatial and temporal priors into disentangled attention pathways. At inference, our Attention Mixing mechanism effectively fuses these learned attentions to produce coherent, persistent, and compositionally consistent 4D reconstructions, even in scenes containing multiple interacting static and dynamic objects. COM4D achieves state-of-the-art performance on both 3D compositional scene reconstruction and 4D dynamic object reconstruction, underscoring the versatility and effectiveness of the approach. Despite these promising results, certain limitations remain. The model’s understanding of motion is learned from data and does not incorporate an explicit physical causality. Consequently, when objects become occluded, it can struggle to reason about their continued trajectory and interaction in a physically plausible manner. Furthermore, COM4D is designed to operate on videos captured from a fixed camera perspective and does not currently support scenarios with camera motion. Future work could explore integrating physical causality to improve reasoning under occlusions and extending it to dynamic camera inputs.
Acknowledgements.
This research was partially funded by the Ministry of Education and Science of Bulgaria (support for INSAIT, part of the Bulgarian National Roadmap for Research Infrastructure). The research was also partly funded by INSAIT-VIVO project on 3D scene understanding and editing.
Supplementary Material
A Training Details
We provide details of the full training losses described in Sec. 3.2. Specifically, eq. (3) provides the equation for $\mathcal{L}_{S}$ and thereafter describes $\mathcal{L}_{T}$ in text. Here, we write $\mathcal{L}_{T}$ formally:
$\mathcal{L}_{T}=\mathbb{E}\left[\sum_{f=1}^{F}\left\|({{}^{f}}\!\boldsymbol{\epsilon}-{{}^{f}}\!\mathbf{z}_{0})-\mathbf{v}_{\theta}({{}^{f}}\!\mathbf{z}_{t_{f}},t_{f},{{}^{f}}\!\mathbf{y})\right\|^{2}\right].$ (4)
The above expression in eq. (4) for the temporal loss is almost identical to the spatial expression in eq. (3). The obvious change is the the frame index $f$ replacing the object index $i$, as each data: $({{}^{f}}\!\mathbf{z}_{0},{{}^{f}}\!\mathbf{y})$ is sampled from DeformingThings [42]. Additionally, the conditional image embeddings ${{}^{f}}\!\mathbf{y}$ are separate for each frame among $F$, unlike in the spatial loss expression, where all $N$ objects use the same image embedding $\mathbf{y}$.
Finally for completeness, we formally write the regularization loss, i.e., the TripoSG [43] loss [^1] The training loss for DiT is not mentioned explicitly in the reference. Please refer to Algorithm 1 in [50] for the rectified flow loss. as follows:
$\mathcal{L}_{R}=\mathbb{E}\left[\left\|(\boldsymbol{\epsilon}-\mathbf{z}_{0})-\mathbf{v}_{\theta}(\mathbf{z}_{t},t,\mathbf{y})\right\|^{2}\right].$ (5)
Note that, each sample in eq. (5) only consists of one object with no temporal evolution. Samples $(\mathbf{z}_{0},\mathbf{y})$ are obtained from the Objaverse training set [15]. Finally, the overall loss is $\mathcal{L}_{S/T/R}$: $\mathcal{L}_{S},\mathcal{L}_{T},\mathcal{L}_{R}$ are sampled with a ratio of $0.35:0.35:0.3$ respectively. The regularization and its sampling ratio of 0.3 is also used by PartCrafter [47] and MIDI [29].
B User Study Details
To quantitatively evaluate our method’s perceptual quality, we conducted a user preference study. The study was administered via Google Forms and compared our full model (with Attention Mixing) against an ablation baseline (without Attention Mixing).
Procedure.
As shown in Fig. 8, participants were presented with a 2D input sequence indicating intended object motion and two corresponding 3D animated samples (labeled (1) and (2)). For each comparison, they answered the question: ”Which sample better matches the input in terms of object placement, motion, and scene structure?” by rating their preference on a 5-point Likert scale (1: Sample 1 is better, 3: Both are about the same, 5: Sample 2 is better). To prevent bias, the assignment of our method to Sample (1) or (2) and the order of the scenes were randomized for each participant.
Figure 8: Our user study interface, run on Google Forms. Participants viewed a 2D input sequence (top) and two 3D results (middle, bottom), then rated their preference on a 5-point scale.
Data Analysis.
All ratings were included in our final analysis. The preference scores reported in the main paper show the complete distribution of judgments across the 5-point Likert scale.
C Evaluation on CMU Panoptic Dataset
To quantitatively assess the temporal consistency and motion accuracy of our model, we performed an evaluation on the CMU Panoptic dataset [32]. This section details our protocol for preparing the data and computing the metrics.
Ground Truth Point Cloud Generation.
We first generated ground truth (GT) point clouds from the raw RGB-D Kinect data provided in the dataset. To ensure a clean and fair comparison, we pre-processed these GT clouds in two steps:
  1. 1.
    Ground Removal: The ground plane was removed using a simple height threshold.
  2. 2.
    Denoising: We applied a statistical outlier removal filter to eliminate stray, floating points in the cloud.
Alignment and Metric Computation.
A key challenge in evaluating generative models is that their outputs are not inherently aligned with the GT coordinate system; they may have an arbitrary scale, rotation, and translation. To address this, we adopted a first-frame alignment protocol.
For each sequence, we independently registered the initial generated mesh (frame 1) from both our full model and the baseline (without Attention Mixing) to the corresponding ground truth point cloud. This one-time alignment transformation (capturing scale, rotation, and translation) was then applied uniformly to all subsequent frames generated by that method for the entire sequence.
Finally, we computed the Chamfer Distance (CD) between our transformed per-frame reconstructions and the GT point clouds. This metric effectively measures how much the predicted motion deviates from the ground truth over time, given an initial registration. A lower accumulated CD indicates a more accurate and temporally consistent motion prediction. Visual examples of these per-frame alignments are shown in Fig. 9, Fig. 10 and Fig. 11. We can observe that not only registered $1^{st}$ frame, but also the remaining frames align well with the GT pointcloud despite the motion, with surprisingly small drift.
Input
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
w/ Mixing
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
w/o Mixing
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 9: Ablation study on mixing components across five time steps in the CMU Panoptic [32] sample. The top row shows the ground truth frames, followed by two views with our mixing strategy and two views without. Gray points denote the ground truth point cloud.
Input
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
w/ Mixing
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
w/o Mixing
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 10: Ablation study on mixing components across five time steps. The top row shows the ground truth frames, followed by two views with our mixing strategy and two views without. Gray points denote the ground truth point cloud.
Input
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
w/ Mixing
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
w/o Mixing
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 11: Ablation study on mixing components across five time steps. The top row shows the ground truth frames, followed by two views with our mixing strategy and two views without. Gray points denote the ground truth point cloud.
D Scalability with Respect to Object Count
COM4D supports denoising with up to 8 parts, and up to 16 parts with additional finetuning. In the static dataset, scenes contain on average 6.19 objects, with the most common configuration being 5 objects (16.67%).
We observe that performance is moderately sensitive to both the number of dynamic objects and the complexity of the static scene. As the number of interacting objects increases, the compositional reasoning task becomes more challenging, leading to gradual degradation in reconstruction quality.
E Additional Qualitative Results on Compositional 4D
Input
Refer to caption Refer to caption Refer to caption
w/ Mixing
Refer to caption Refer to caption Refer to caption
w/o Mixing
Refer to caption Refer to caption Refer to caption
Input
Refer to caption Refer to caption Refer to caption
w/ Mixing
Refer to caption Refer to caption Refer to caption
w/o Mixing
Refer to caption Refer to caption Refer to caption
Figure 12: Ablation study on mixing components for various sequences. For each sample, we show input frames, results with our mixing strategy, and results without. In particular, the chair pose and the interaction between the dynamics (the lady shaking hands with the man) and the dynamic and static (lady and the chair) are captured incorrectly without mixing.
F Additional Qualitative Results with Moving Camera
COM4D assumes a fixed camera when reconstructing compositional scenes containing both static and dynamic objects, as camera motion is inherently entangled with static scene geometry. In such settings, disentangling camera motion from scene structure is fundamentally ambiguous without additional constraints or supervision.
That said, COM4D is not inherently restricted to static-camera scenarios. In cases where the scene consists only of dynamic objects, the model can still recover consistent relative 4D structure and motion over time. This is because COM4D relies on learning relative spatial relationships and temporal coherence, rather than absolute camera-referenced geometry.
Input
Refer to caption Refer to caption Refer to caption Refer to caption
Reconstr.
Refer to caption Refer to caption Refer to caption Refer to caption
Novel View
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 13: Qualitative results for a sample with dynamic camera.
G Physical Plausibility and Object Interactions
COM4D is a fully data-driven framework and does not explicitly enforce physical constraints such as contact or collision avoidance. Instead, interactions between multiple objects are modeled implicitly through attention mechanisms in the latent space.
To assess whether this implicit modeling yields physically plausible reconstructions, we evaluate mesh interpenetration across 8 video sequences by measuring the intersection-over-union (IoU) between reconstructed object meshes. We observe a low average overlap of $\text{IoU}=0.0096$, indicating minimal interpenetration between interacting objects.
For reference, we report an IoU of $0.0018$ on 3D-FRONT, where ground-truth scenes exhibit little to no physical contact between objects (see Tab. 2). The small gap between these values suggests that COM4D maintains a comparable level of physical plausibility despite not explicitly modeling physical constraints.
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 14: Qualitative results on interaction scenarios. Top row shows input frames, and bottom shows 4D reconstructions.
H Comparison with Optimization-Based Baselines
We compare COM4D against the recent optimization-based method Shape of Motion (SOM) [80]. SOM struggles to recover complete geometry under occlusions, often producing incomplete or noisy reconstructions (see Fig. 15).
Quantitatively, SOM achieves a Chamfer Distance (CD) of 11.30 cm, whereas COM4D achieves 7.42 cm on [32, 33]. This gap highlights the advantage of our generative formulation, which leverages learned spatial and temporal priors to produce more complete and coherent reconstructions in challenging compositional settings.
                       Input
Refer to caption Refer to caption Refer to caption Refer to caption
                  SOM [80]
Refer to caption Refer to caption Refer to caption Refer to caption
                   Ours
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 15: Qualitative comparison across methods.
I Evaluation Using a Synthetic Dataset
We additionally evaluate our method on a synthetic 4D dataset since a suitable real-captured 4D scene benchmark is unavailable. We first generate candidate synthetic scenes by composing one or more dynamic humanoid motion sequences with several static objects. Each dynamic object is selected from a motion sequence, sub-sampled with a fixed frame stride, and placed into a shared normalized world coordinate system. Static assets are then scaled, rotated around the vertical axis, and translated so that they fit around the dynamic trajectories without excessive overlap. For every time step, we export a combined GLB that contains the static geometry and the current pose of each dynamic object, and we also save metadata describing the source assets, transforms, per-frame bounds, and global centering applied to the scene.
For visualization, each frame is rendered from a fixed orthographic camera with consistent framing across the entire sequence. The renderer keeps the scene centered and uses the same camera and lighting conventions for all frames within a sample, which makes frame-to-frame motion directly comparable.
For quantitative evaluation, we compare the generated scene-level mesh against the static part of the reference scene and compute dynamic scores per moving object and per frame. For static geometry, we compute one CD/F-score pair per scene and report the arithmetic mean over scenes. For dynamic content, we first average all valid dynamic per-frame scores within each sample, and then average these per-sample dynamic means across the whole set. This prevents samples with more moving objects or more valid frames from dominating the final measure.
Component CD $\downarrow$ F-score $\uparrow$
Static scene geometry 0.293 0.418
Dynamic objects 0.284 0.434
Table 4: Quantitative results on the synthetic 4D evaluation set. Lower CD is better; higher F-score is better.
These synthetic-scene scores are worse than the corresponding single-object 4D and static compositional 3D scene results. The main reason is that, although the reconstructed objects themselves and their coarse arrangements are often visually consistent with the input, the relative positioning of objects is still somewhat inaccurate. Small errors in placement and pairwise spatial relationships accumulate at the full-scene level and are directly penalized by both CD and F-score, leading to lower quantitative performance than in simpler single-object or static-only settings.
J Additional Qualitative Results on Synthetic Compositional 4D Dataset
In Fig. 16, Fig. 17, Fig. 18 we show additional qualitative results on our synthetic compositional 4D dataset.
Input Refer to caption Refer to caption Refer to caption Refer to caption
COM4D Refer to caption Refer to caption Refer to caption Refer to caption
Input Refer to caption Refer to caption Refer to caption Refer to caption
COM4D Refer to caption Refer to caption Refer to caption Refer to caption
Input Refer to caption Refer to caption Refer to caption Refer to caption
COM4D Refer to caption Refer to caption Refer to caption Refer to caption
Input Refer to caption Refer to caption Refer to caption Refer to caption
COM4D Refer to caption Refer to caption Refer to caption Refer to caption
Figure 16: Additional qualitative results on the synthetic compositional 4D dataset.
Input Refer to caption Refer to caption Refer to caption Refer to caption
COM4D Refer to caption Refer to caption Refer to caption Refer to caption
Input Refer to caption Refer to caption Refer to caption Refer to caption
COM4D Refer to caption Refer to caption Refer to caption Refer to caption
Input Refer to caption Refer to caption Refer to caption Refer to caption
COM4D Refer to caption Refer to caption Refer to caption Refer to caption
Input Refer to caption Refer to caption Refer to caption Refer to caption
COM4D Refer to caption Refer to caption Refer to caption Refer to caption
Figure 17: Additional qualitative results on the synthetic compositional 4D dataset.
Input Refer to caption Refer to caption Refer to caption Refer to caption
COM4D Refer to caption Refer to caption Refer to caption Refer to caption
Input Refer to caption Refer to caption Refer to caption Refer to caption
COM4D Refer to caption Refer to caption Refer to caption Refer to caption
Input Refer to caption Refer to caption Refer to caption Refer to caption
COM4D Refer to caption Refer to caption Refer to caption Refer to caption
Input Refer to caption Refer to caption Refer to caption Refer to caption
COM4D Refer to caption Refer to caption Refer to caption Refer to caption
Figure 18: Additional qualitative results on the synthetic compositional 4D dataset.
K Additional Qualitative Results on 4D Object Reconstruction
In  Fig. 19,  Fig. 20 and  Fig. 21 we show more qualitative results of our model and baselines on single object 4D object reconstruction.
Input / GT Ours TripoSG V2M4 GVFD L4GM
Input $\angle 0^{\circ}$ Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
render $\angle 90^{\circ}$ Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Input $\angle 0^{\circ}$ Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
render $\angle 90^{\circ}$ Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Input $\angle 0^{\circ}$ Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
render $\angle 90^{\circ}$ Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Input $\angle 0^{\circ}$ Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
render $\angle 90^{\circ}$ Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Input $\angle 0^{\circ}$ Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
render $\angle 90^{\circ}$ Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Input $\angle 0^{\circ}$ Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
render $\angle 90^{\circ}$ Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 19: Qualitative 4D generation comparisons from Objaverse [15] showing three subjects at two time steps. For each model, we show the reconstructed input view (top) and a rendered novel view (bottom). We display the ground truth novel view in the bottom left.
Input / GT Ours TripoSG V2M4 GVFD L4GM
Input $\angle 0^{\circ}$ Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
render $\angle 90^{\circ}$ Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Input $\angle 0^{\circ}$ Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
render $\angle 90^{\circ}$ Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Input $\angle 0^{\circ}$ Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
render $\angle 90^{\circ}$ Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Input $\angle 0^{\circ}$ Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
render $\angle 90^{\circ}$ Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Input $\angle 0^{\circ}$ Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
render $\angle 90^{\circ}$ Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Input $\angle 0^{\circ}$ Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
render $\angle 90^{\circ}$ Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 20: Further qualitative 4D generation comparisons from Objaverse [15] showing two subjects of Objaverse [15], at three time steps. For each model, we show the reconstructed input view (top) and a rendered novel view (bottom). We display the ground truth novel view in the bottom left. The novel view in particular highlights the discrepancies of the methods’ outputs from the ground truth. Both TripoSG and V2M4 show moderate and consistent misalignment. Similar misalignment, particularly in rotation and skeletal pose is apparent in GVFD, while L4GM often fails to provide good novel views. Typically, our method shows far less shape or pose misalignment, as reflected in the quantitative metrics.
Input / GT Ours TripoSG GVFD L4GM
Input $\angle 0^{\circ}$ Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
render $\angle 90^{\circ}$ Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Input $\angle 0^{\circ}$ Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
render $\angle 90^{\circ}$ Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Input $\angle 0^{\circ}$ Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
render $\angle 90^{\circ}$ Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Input $\angle 0^{\circ}$ Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
render $\angle 90^{\circ}$ Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Input $\angle 0^{\circ}$ Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
render $\angle 90^{\circ}$ Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Input $\angle 0^{\circ}$ Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
render $\angle 90^{\circ}$ Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 21: Qualitative 4D generation comparisons from DeformingThings [42] on two subjects at three time steps. For each model, we show the reconstructed input view (top) and a rendered novel view (bottom). We display the ground truth novel view in the bottom left. Compared to Fig. 20, the sequences contain stronger motion thus showing even larger difference in performances of our method. In particular, V2M4 fails completely due to the large motion.
L Additional Qualitative Results on 3D Scene Generation
Input Ours PartCrafter [47] MIDI [29]
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 22: Qualitative comparison across ours, PartCrafter [47] and MIDI [29]. We show qualitatively how our method shows better performance, as shown by the quantitative metrics. Largely, this is due to the consistent reconstruction of all parts and their accurate composition. Both PartCrafter [47] and MIDI [29] often miss large objects, e.g., the bed.